# Constraint on initial conditions of one-dimensional expanding fluids from nonlinear causality

### v1

@@ III.3 Constraint on Initial Conditions
Any solutions that pass through the acausal regions are by no means acceptable from a viewpoint of relativistic theory. The initial values of the shear pressure $\phi_{0}$, the energy density $e_{0}$, and the resulting pressure $p_{0}$ calculated from EoS must be chosen so that the solutions never intrude the acausal regions. Thus the initial conditions are considerably constrained in $\tau_{0}/\tau_{\pi 0}$-$\phi_{0}/(e_{0}+p_{0})$ plane from causality argument as shown in Fig. 3. Wherever an initial value is taken from the region of acausal initial conditions in Fig. (3), the solution of Eqs. (9) and (10) definitely violates causality at least during its evolution. On the other hand, wherever an initial value is taken from the region of causal initial conditions in Fig. (3), the solution is always guaranteed to be causal during its evolution.
Figure 3: The regions of acausal (red) and causal (blue) initial conditions in $\tau_{0}/\tau_{\pi 0}$-$\phi_{0}/(e_{0}+p_{0})$ plane. See the text for details.
It would be instructive to draw the regions of initial conditions for their solutions to be causal or acausal in another plane. For a fixed initial proper time $\tau_{0}=0.2$, $0.8$, or $2.0$ fm, the regions of acausal and causal initial conditions are shown in Fig. 4. For an extremely early initial proper time ($\tau_{0}=0.2$ fm), there is no room for solutions to be guaranteed to be causal as shown in Fig. 4 (a). Even if the fluids start from local equilibrium state, the expansion scalar in one-dimensional expanding system, $\partial_{\mu}u_{\mathrm{Bj}}^{\mu}=1/\tau$, becomes tremendously large for extremely early initial proper time and, consequently, fluids are quickly far away from local equilibrium state. As $\tau_{0}$ increases to 0.8 (Fig. 4 (b)) and 2.0 fm (Fig. 4 (c)), the region of causal initial conditions appears more likely.
Figure 4: The regions of acausal (red) and causal (blue) initial conditions in $e_{0}$-$\phi_{0}$ plane at the initial time, (a) $\tau_{0}=0.2$, (b) $0.8$, and (c) $2.0$ fm. See the text for details.
@@
where $S$ is an effective area of the overlap region in the transverse plane. From Eq. (33), one obtains $e_{0}\tau_{0}\sim 5$ GeV/fm2 in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV Adams et al. (2004) and $\sim 14$ GeV/fm2 in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV Chatrchyan et al. (2012). The regions of acausal and causal initial conditions with fixing $e_{0}\tau_{0}$ are shown in Fig. 5. One finds from Fig. 5 that the acceptable region for the system guaranteed to be causal is highly constrained from experimental inputs. A point at the maximum energy density in the region guaranteed to be causal in Fig. 5 corresponds to the minimum initial proper time, $\tau_{0}/\tau_{\pi}$, in Fig. 3. Therefore the minimum initial time is obtained from a combination of nonlinear causality and the measured transverse energy.
@@
By assuming three flavor massless ideal gas of the QGP ($N_{f}=3$ and $d_{\mathrm{eff}}=47.5$), the energy density can be transformed into the temperature:
@@
$\frac{\tau_{0,\mathrm{min}}}{\tau_{\pi}}=\frac{\tau_{0,\mathrm{min}}T_{0,\mathrm{max}}}{C_{\tau_{\pi}}}\sim 5.3.$ (36)
@@
From Eqs. (33) and (37), maximum energy density and minimum initial proper time are $e_{0,\mathrm{max}}\sim 5$ GeV/fm3 and $\tau_{0,\mathrm{min}}\sim 1$ fm in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and $e_{0,\mathrm{max}}\sim 20$ GeV/fm3 and $\tau_{0,\mathrm{min}}\sim 0.7$ fm in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV. If one relies strictly on the necessary and the sufficient conditions from nonlinear causality in one-dimensional expansion, these pose the minimum bound of the initial proper time. However, if one imposes the necessary conditions only, the broader regions are available for the initial conditions in one-dimensional expansion of the conformal fluids and the earlier initial time could be acceptable.
@@ III.4 Results with lattice EoS
For a more realistic analysis, results with the conformal EoS are compared with the ones with the lattice EoS in this subsection. Note that we replace the model EoS only and keep the equation of motion and the transport coefficients as in the previous subsections.
Since the system is no longer conformal one, the solutions of Eqs. (9) and (10) depend on the energy scale such as initial temperature or initial energy density. Thus, we choose $e_{0}=5$ and $14$ GeV/fm3 at RHIC and LHC energies, respectively, as the typical energy densities. From the causality conditions for these two energy scales, we obtain the inequalities among thermodynamic variables and the shear pressure as follows: In the case of $e_{0}=5$ GeV/fm3,
@@
Figure 6: The same as Fig. 3 except for employing lattice EoS. We set (a) $e_{0}=5$ GeV/fm3 and (b) $e_{0}=14$ GeV/fm3.
@@
One finds in Fig. 7 that maximum initial energy density and minimum initial proper time are $e_{0,\mathrm{max}}\sim 11$ GeV/fm3 and $\tau_{0,\mathrm{min}}\sim 0.5$ fm in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and $e_{0,\mathrm{max}}\sim 35$ GeV/fm3 and $\tau_{0,\mathrm{min}}\sim 0.4$ fm in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV. In comparison with the results from the conformal EoS, it turns out that the acceptable regions are broadened in the lattice EoS case.
@@
Figure 7: The same as Fig. 5 except for employing lattice EoS. (a) $e_{0}\tau_{0}=5$ GeV/fm2 at RHIC and (b) $14$ GeV/fm2 at LHC energy.
@@ IV Summary
We first obtained the necessary and the sufficient conditions of nonlinear causality in one-dimensional expanding systems which obey the BRSSS equation. The necessary and the sufficient conditions are stated as 9 and 8 inequalities, respectively, among the thermodynamic variables, the shear pressure, and the transport coefficients. Using the conformal EoS and a particular set of transport coefficients, we obtained $-0.47<\phi/(e+p)<0.23$ and $-0.07<\phi/(e+p)<0.07$ from the necessary and the sufficient conditions, respectively. From these results, we concluded that the inverse Reynolds number must be small, meaning that the system must be close to the local equilibrium, so as to obey the relativistic causality.
@@
With additional constraints from measured transverse energy per rapidity at RHIC and LHC, we further constrained the regions of acausal and causal initial conditions. We found the acceptable initial conditions for the system guaranteed to be causal are highly constrained from a combination of nonlinear causality and the measured transverse energy. We derived a simple formula to quantify minimum initial time as $\tau_{0,\mathrm{min}}\enskip(\mathrm{fm})\approx 1.6\times[(1/S)dE_{T}/dy\enskip(\mathrm{GeV/fm}^{2})]^{-1/3}$ for $N_{f}=3$ conformal system. Using this formula, maximum initial energy density and minimum initial proper time are obtained as $e_{0,\mathrm{max}}\sim 5$ GeV/fm3 and $\tau_{0,\mathrm{min}}\sim 1$ fm in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and $e_{0,\mathrm{max}}\sim 20$ GeV/fm3 and $\tau_{0,\mathrm{min}}\sim 0.7$ fm in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV in the conformal limit. Although these constraints were obtained from the necessary and the sufficient conditions of nonlinear causality, these could be relaxed if one chooses only the necessary conditions.
@@
By employing the lattice EoS, we refined the results obtained using the conformal EoS. Maximum initial energy density and minimum initial proper time are obtained as $e_{0,\mathrm{max}}\sim 11$ GeV/fm3 and $\tau_{0,\mathrm{min}}\sim 0.5$ fm in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and $e_{0,\mathrm{max}}\sim 35$ GeV/fm3 and $\tau_{0,\mathrm{min}}\sim 0.4$ fm in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV. In comparison with the results with the conformal EoS case, a bigger $e_{0,\mathrm{max}}$ and a smaller $\tau_{0,\mathrm{min}}$ are acceptable in the case of lattice EoS.
@@
In this paper, we employed the BRSSS equation as a constitutive equation to demonstrate how the inverse Reynolds number is related with violation of causality. Since the constitutive relation can be also obtained in various approaches, varieties of the second order constitutive equations can be found in the literature (e.g., Refs. Hiscock and Lindblom (1983); Baier et al. (2006); Betz et al. (2009a, b); Denicol et al. (2012); Monnai and Hirano (2010); Tsumura et al. (2007); Tsumura and Kunihiro (2010, 2013); Jaiswal (2013)). This could give a theoretical systematic uncertainty for constraints on initial conditions. A detailed analysis by employing other constitutive equations as well as parameter dependence of causality conditions on transport coefficients will be made in future analysis.
@@
We emphasize here that we are not arguing against the existence of attractor behavior in this paper. Rather, we claim the framework other than the Müller–Israel–Stewart type constitutive equations in relativistic dissipative dynamics is indispensable toward comprehensive understanding of hydrodynamization and/or thermalization in relativistic heavy-ion collisions. A microscopic description based on the kinetic theory might be one of them Heller et al. (2018); Blaizot and Yan (2018); Romatschke (2018); Behtash et al. (2018); Heller and Svensson (2018); Strickland (2018); Denicol and Noronha (2020); Heller et al. (2020); Kamata et al. (2020); Almaalol et al. (2020); Chattopadhyay et al. (2022); Jaiswal et al. (2022); Brewer et al. (2024). Just as we have examined the conditions for the applicability of relativistic hydrodynamics through nonlinear causality in this paper, we need to carefully consider the applicability of the Boltzmann equation. In particular, the density of the system must be sufficiently low so that the truncation of the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy is justified Lifshitz and Pitaevskii (1981). This would be challenging in the very early stage of high-energy nuclear collisions in which the parton density is extremely high.
@@
The validity of numerical simulations of relativistic dissipative hydrodynamics has been scrutinized through the necessary and the sufficient conditions of nonlinear causality Plumberg et al. (2022); Chiu and Shen (2021); da Silva et al. (2023); Krupczak et al. (2024); Domingues et al. (2024) and it turns out that causality is indeed violated in the early stage and/or in peripheral regions. The violation of causality is intimately related with instability (see, e.g., Ref. Gavassino (2022)). Nevertheless, numerical simulations of relativistic dissipative hydrodynamics seemingly work well without any instability behaviors. However, it is by no means trivial whether instabilities arise when hydrodynamic fluctuations act as perturbations on a local equilibrium system. This becomes an important issue in sophisticated (3+1)-dimensional numerical simulations based on relativistic fluctuating hydrodynamics Murase (2015); Sakai et al. (2020, 2022); Kuroki et al. (2023).
Acknowledgements
The work by T.H. was partly supported by JSPS KAKENHI Grant Number JP23K03395.
Appendix A One-dimensional expansion of conformal system
In this Appendix A, we discuss “time” evolution of conformal system which undergoes one-dimensional boost invariant expansion following the discussion in Refs. Heller and Spalinski (2015); Florkowski et al. (2018). In particular, we exhibit gradient vectors of solutions to capture the global behavior of them.
@@ Appendix A One-dimensional expansion of conformal system
Here $\tau$ and $T$ are the proper time and the temperature, respectively. Note that our definition of a variable $f$ is a factor $3/2$ different from the original one Heller and Spalinski (2015) for later purpose. By using the equation of motion (9) together with $e\propto T^{4}$ and $p=e/3$ in the relativistic conformal system, a variable $f$ can be also written by using the inverse Reynolds number (12) as
@@
Figure 8: Gradient of solutions for Eq. (48) at various points in $w$$f$ plane. Each arrow shows a vector ($\Delta w$, $\Delta f$) and the upper and the lower curves show the solution of $df/dw=0$, $f=f_{0}^{+}(w)$ and $f=f_{0}^{-}(w)$, respectively.
@@
where $f_{0}^{\pm}(w\rightarrow 0)=1\pm\frac{1}{2}\sqrt{\frac{C_{\eta}}{C_{\tau_{\pi}}}}$. Figure 8 shows “flow” vectors $(\Delta w,\Delta f)$ and $f=f_{0}^{\pm}(w)$ in $w$-$f$ plane. From Fig. 8, $df/dw>0$ in the region between $f=f_{0}^{+}(w)$ and $f=f_{0}^{-}(w)$, while $df/dw<0$ in the others. Gradient of a solution changes its sign when it hits $f=f_{0}^{+}(w)$. It should be noted that $f=f_{0}^{+}(w)$ is not the hydrodynamic attractor solution but merely the boundary of changing the sign of $df/dw$. The other solution, $f=f_{0}^{-}(w)$, is “unstable” in the sense that the solutions, $f(w)$, deviate from $f=f_{0}^{-}(w)$ due to $df/dw>0$ ($df/dw<0$) above (below) this boundary. It is often denoted that any initial conditions converge to the so-called hydrodynamic attractor in this system. However, if the system starts from an initial condition below $f=f_{0}^{-}(w)$, the solution no longer converges to the hydrodynamic attractor.

### v2

@@ III.3 Constraint on Initial Conditions
Any solutions that pass through the acausal regions are by no means acceptable from a viewpoint of relativistic theory. The initial values of the shear pressure $\phi_{0}$, the energy density $e_{0}$, and the resulting pressure $p_{0}$ calculated from the EoS must be chosen so that the solutions never intrude the acausal regions. Thus the initial conditions are considerably constrained in the $\tau_{0}/\tau_{\pi 0}$-$\phi_{0}/(e_{0}+p_{0})$ plane from the causality argument as shown in Fig. 3. Wherever an initial value is taken from the region of acausal initial conditions in Fig. 3, the solution of Eqs. (9) and (10) definitely violates causality at least during its evolution. On the other hand, wherever an initial value is taken from the region of causal initial conditions in Fig. 3, the solution is always guaranteed to be causal during its evolution.
Figure 3: The regions of acausal (red) and causal (blue) initial conditions in the $\tau_{0}/\tau_{\pi 0}$-$\phi_{0}/(e_{0}+p_{0})$ plane. See the text for details.
It would be instructive to draw the regions of initial conditions whose solutions are causal or acausal in another plane. For a fixed initial proper time $\tau_{0}=0.2$, $0.8$, or $2.0$ fm, the regions of acausal and causal initial conditions are shown in Fig. 4. For an extremely early initial proper time ($\tau_{0}=0.2$ fm), there is no room for solutions that are guaranteed to be causal, as shown in Fig. 4 (a). Even if the fluids start from a local equilibrium state, the expansion scalar in a one-dimensional expanding system, $\partial_{\mu}u_{\mathrm{Bj}}^{\mu}=1/\tau$, becomes tremendously large for extremely early initial proper time and, consequently, fluids are quickly far away from the local equilibrium state. As $\tau_{0}$ increases to 0.8 [Fig. 4 (b)] and 2.0 fm [Fig. 4 (c)], the region of causal initial conditions expands.
Figure 4: The regions of acausal (red) and causal (blue) initial conditions in the $e_{0}$-$\phi_{0}$ plane at the initial time, (a) $\tau_{0}=0.2$, (b) $0.8$, and (c) $2.0$ fm. See the text for details.
@@
where $S$ is an effective area of the overlap region in the transverse plane. From Eq. (33), one obtains $e_{0}\tau_{0}\approx 5$ GeV/fm2 in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV Adams et al. (2004) and $\approx 14$ GeV/fm2 in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV Chatrchyan et al. (2012). The regions of acausal and causal initial conditions with fixed $e_{0}\tau_{0}$ are shown in Fig. 5. One finds from Fig. 5 that the acceptable region for a system guaranteed to be causal is highly constrained from experimental inputs. A point at the maximum energy density in the region guaranteed to be causal in Fig. 5 corresponds to the minimum initial proper time, $\tau_{0}/\tau_{\pi}$, in Fig. 3. Therefore the minimum initial time is obtained from a combination of nonlinear causality and the measured transverse energy.
@@
By assuming a three-flavor massless ideal gas for the QGP ($N_{f}=3$ and $d_{\mathrm{eff}}=47.5$), the energy density can be transformed into the temperature:
@@
$\frac{\tau_{0,\mathrm{min}}}{\tau_{\pi}}=\frac{\tau_{0,\mathrm{min}}T_{0,\mathrm{max}}}{C_{\tau_{\pi}}}\approx 5.3.$ (36)
@@
From Eqs. (33) and (37), maximum energy density and minimum initial proper time are $e_{0,\mathrm{max}}\approx 5$ GeV/fm3 and $\tau_{0,\mathrm{min}}\approx 1$ fm in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and $e_{0,\mathrm{max}}\approx 20$ GeV/fm3 and $\tau_{0,\mathrm{min}}\approx 0.7$ fm in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV. If one relies strictly on the sufficient conditions from nonlinear causality in one-dimensional expansion, these impose the minimum bound of the initial proper time. However, if one imposes the necessary conditions only, the broader regions are available for the initial conditions in one-dimensional expansion of the conformal fluids and the earlier initial time could be acceptable.
@@ III.4 Results with lattice EoS
For a more realistic analysis, results with the conformal EoS are compared in this subsection with the ones using the lattice EoS. Note that we replace the model EoS only and keep the equation of motion and the transport coefficients as in the previous subsections.
Since the system is no longer a conformal one, the solutions of Eqs. (9) and (10) depend on the energy scale such as initial temperature or initial energy density. Thus, we choose $e_{0}=5$ and $14$ GeV/fm3 at RHIC and LHC energies, respectively, as the typical energy densities. From the causality conditions for these two energy scales, we obtain the inequalities among thermodynamic variables and the shear pressure as follows: In the case of $e_{0}=5$ GeV/fm3,
@@
Figure 6: The same as Fig. 3 but employing the lattice EoS. We set (a) $e_{0}=5$ GeV/fm3 and (b) $e_{0}=14$ GeV/fm3.
@@
One finds in Fig. 7 that maximum initial energy density and minimum initial proper time are $e_{0,\mathrm{max}}\approx 11$ GeV/fm3 and $\tau_{0,\mathrm{min}}\approx 0.5$ fm in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and $e_{0,\mathrm{max}}\approx 35$ GeV/fm3 and $\tau_{0,\mathrm{min}}\approx 0.4$ fm in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV. In comparison with the results from the conformal EoS, it turns out that the acceptable regions are broadened in the lattice EoS case.
@@
Figure 7: The same as Fig. 5 but employing the lattice EoS. (a) $e_{0}\tau_{0}=5$ GeV/fm2 at RHIC and (b) $14$ GeV/fm2 at LHC energy.
@@ IV Summary
We first obtained the necessary and the sufficient conditions of nonlinear causality in one-dimensional expanding systems which obey the BRSSS equation. The necessary and the sufficient conditions are stated as nine and eight inequalities, respectively, among the thermodynamic variables, the shear pressure, and the transport coefficients. Using the conformal EoS and a particular set of transport coefficients, we obtained $-0.47<\phi/(e+p)<0.23$ and $-0.07<\phi/(e+p)<0.07$ from the necessary and the sufficient conditions, respectively. From these results, we concluded that the inverse Reynolds number must be small, meaning that the system must be close to the local equilibrium, in order to obey the relativistic causality.
@@
With additional constraints from measured transverse energy per rapidity at RHIC and LHC, we further constrained the regions of acausal and causal initial conditions. We found the acceptable initial conditions for the systems guaranteed to be causal are highly constrained from a combination of nonlinear causality and the measured transverse energy. We derived a simple formula to quantify minimum initial time as $\tau_{0,\mathrm{min}}\enskip(\mathrm{fm})\approx 1.6\times[(1/S)dE_{T}/dy\enskip(\mathrm{GeV/fm}^{2})]^{-1/3}$ for an $N_{f}=3$ conformal system. Using this formula, maximum initial energy density and minimum initial proper time are obtained as $e_{0,\mathrm{max}}\approx 5$ GeV/fm3 and $\tau_{0,\mathrm{min}}\approx 1$ fm in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and $e_{0,\mathrm{max}}\approx 20$ GeV/fm3 and $\tau_{0,\mathrm{min}}\approx 0.7$ fm in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV in the conformal limit. Although these constraints were obtained from the sufficient conditions of nonlinear causality, these could be relaxed if one chooses only the necessary conditions.
@@
By employing the lattice EoS, we refined the results obtained using the conformal EoS. Maximum initial energy density and minimum initial proper time are obtained as $e_{0,\mathrm{max}}\approx 11$ GeV/fm3 and $\tau_{0,\mathrm{min}}\approx 0.5$ fm in top 5% Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and $e_{0,\mathrm{max}}\approx 35$ GeV/fm3 and $\tau_{0,\mathrm{min}}\approx 0.4$ fm in top 5% Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV. In comparison with the results with the conformal EoS case, a bigger $e_{0,\mathrm{max}}$ and a smaller $\tau_{0,\mathrm{min}}$ are acceptable in the case of lattice EoS.
@@
In this paper, we employed the BRSSS equation as a constitutive equation to demonstrate how the inverse Reynolds number is related to violation of causality. Since the constitutive relation can be also obtained in various approaches, varieties of the second order constitutive equations can be found in the literature (e.g., Refs. Hiscock and Lindblom (1983); Baier et al. (2006); Betz et al. (2009a, b); Denicol et al. (2012); Monnai and Hirano (2010); Tsumura et al. (2007); Tsumura and Kunihiro (2010, 2013); Jaiswal (2013)). This could give a theoretical systematic uncertainty for constraints on initial conditions. A detailed analysis by employing other constitutive equations as well as parameter dependence of causality conditions on transport coefficients will be made in future analysis.
@@
We emphasize here that we are not arguing against the existence of attractor behavior in this paper. Rather, we claim that a framework other than the Müller–Israel–Stewart type constitutive equations in relativistic dissipative dynamics is indispensable toward comprehensive understanding of hydrodynamization and/or thermalization in relativistic heavy-ion collisions. A microscopic description based on the kinetic theory might be such a framework Heller et al. (2018); Blaizot and Yan (2018); Romatschke (2018); Behtash et al. (2018); Heller and Svensson (2018); Strickland (2018); Denicol and Noronha (2020); Heller et al. (2020); Kamata et al. (2020); Almaalol et al. (2020); Chattopadhyay et al. (2022); Jaiswal et al. (2022); Brewer et al. (2024). Just as we have examined the conditions for the applicability of relativistic hydrodynamics through nonlinear causality in this paper, we need to carefully consider the applicability of the Boltzmann equation. In particular, the density of the system must be sufficiently low so that the truncation of the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy is justified Lifshitz and Pitaevskii (1981). This would be challenging in the very early stage of high-energy nuclear collisions in which the parton density is extremely high.
@@
The validity of numerical simulations of relativistic dissipative hydrodynamics has been scrutinized through the necessary and the sufficient conditions of nonlinear causality Plumberg et al. (2022); Chiu and Shen (2021); da Silva et al. (2023); Krupczak et al. (2024); Domingues et al. (2024) and it turns out that causality is indeed violated in the early stage and/or in peripheral regions. The violation of causality is intimately related to instability (see, e.g., Ref. Gavassino (2022)). Nevertheless, numerical simulations of relativistic dissipative hydrodynamics seemingly work well without any instability behaviors. However, it is by no means trivial whether instabilities arise when hydrodynamic fluctuations act as perturbations on a local equilibrium system. This becomes an important issue in sophisticated (3+1)-dimensional numerical simulations based on relativistic fluctuating hydrodynamics Murase (2015); Sakai et al. (2020, 2022); Kuroki et al. (2023).
Acknowledgement
The work by T.H. was partly supported by JSPS KAKENHI Grant No. JP23K03395.
Appendix A One-dimensional expansion of a conformal system
In this Appendix A, we discuss “time” evolution of a conformal system which undergoes one-dimensional boost invariant expansion following the discussion in Refs. Heller and Spalinski (2015); Florkowski et al. (2018). In particular, we present gradient vectors of solutions to capture those solutions’ global behavior.
@@ Appendix A One-dimensional expansion of a conformal system
Here $\tau$ and $T$ are the proper time and the temperature, respectively. Note that our definition of a variable $f$ is a factor $3/2$ different from the original one Heller and Spalinski (2015) for later purposes. By using the equation of motion (9) together with $e\propto T^{4}$ and $p=e/3$ in the relativistic conformal system, a variable $f$ can be also written by using the inverse Reynolds number (12) as
@@
Figure 8: Gradient of solutions for Eq. (48) at various points in the $w$$f$ plane. Each arrow shows a vector ($\Delta w$, $\Delta f$) and the upper and the lower curves show the solution of $df/dw=0$, for $f=f_{0}^{+}(w)$ and $f=f_{0}^{-}(w)$, respectively.
@@
where $f_{0}^{\pm}(w\rightarrow 0)=1\pm\frac{1}{2}\sqrt{\frac{C_{\eta}}{C_{\tau_{\pi}}}}$. Figure 8 shows “flow” vectors $(\Delta w,\Delta f)$ and $f=f_{0}^{\pm}(w)$ in the $w$-$f$ plane. From Fig. 8, $df/dw>0$ in the region between $f=f_{0}^{+}(w)$ and $f=f_{0}^{-}(w)$, while $df/dw<0$ in the others. The gradient of a solution changes its sign when it hits $f=f_{0}^{+}(w)$. It should be noted that $f=f_{0}^{+}(w)$ is not the hydrodynamic attractor solution but merely the boundary of changing the sign of $df/dw$. The other solution, $f=f_{0}^{-}(w)$, is “unstable” in the sense that the solutions, $f(w)$, deviate from $f=f_{0}^{-}(w)$ due to $df/dw>0$ ($df/dw<0$) above (below) this boundary. It is often stated that any initial conditions converge to the so-called hydrodynamic attractor in this system. However, if the system starts from an initial condition below $f=f_{0}^{-}(w)$, the solution no longer converges to the hydrodynamic attractor.